OpenStudy (anonymous):

Stable High School has a total of 112 boys and girls who play sports. If the number of girls, g is 16 more than twice the number of boys, b, how many girls play sports at this high school? Which system of equations would help you solve this problem? Answer b + g = 112 g = 2b + 16 b + g = 112 b = 2g + 16 None of the systems above would solve this problem. b - g = 112 b = 2g - 16

OpenStudy (anonymous):

hi there, so, boys are B and girls are G. one thing you know is that the total number of students is 112, so, how would you translate that into an equation?

OpenStudy (anonymous):

im not sure.... :(

OpenStudy (anonymous):

The sum of two things is just A+B

OpenStudy (anonymous):

would it be the second one

OpenStudy (anonymous):

well forget the multiple choice for a second, let's just develop the equations on our own

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

b+g=112 g=2b+16

OpenStudy (anonymous):

the sum of two things is A+B, and the equals sign means that two things have the same value. For example, "C=D" means that C and D are the same. So, the number of boys is B and the number of girls is G, what's the sum of boys and girls?

OpenStudy (anonymous):

\[\text{Solve}[\{b+g=112,g=16+2b\},\{b,g\}]\]\[\{b\to 32,g\to 80\} \]